in the past lessons,you had collected real data set to be described using statistical values such as the type of the breakfast your relative have.using the same or similar dataset,formulate a problem which you will solve using the normal curve concept .discuss your procedure in doing this task


1
Expert's answer
2021-05-31T18:20:51-0400

For example, we dataset {"x_i"} of different prices for breakfast, which are normally distributed.

And our task is find probability that breakfast costs less than some value X.

Then we calculate mean and standard deviation of prices:

mean:

"\mu=\frac{\sum x_i}{N}"

standard deviation:

"\sigma=\sqrt{\frac{\sum(x_i-\mu)^2}{M}}"

calculate z-value:

"Z=\frac{X-\mu}{\sigma}"

from z-values table take probability P(z<Z)

and find probability for price:

"P(x<X)=P(z<Z)"